This is a pre-algebra question and I'm reviewing a chapter that covers equations and applications involving fractions and I've come across this problem which is very puzzling. I know it has to be a simple answer but it is not coming to me.
12m=8
The answer is 2/3
I can't get this answer no matter what I do. Could you help me I"m feeing pretty lame.

Answer by richwmiller(17219)

2/3 is the reduced fraction
12m=8
m=8/12
reduce the fraction
factor the numerator and denominator
m=2*4/3*4
m=2/3